OpenStudy (anonymous):

find a.b if |a|=2,|b|=5 and the angle between a and b is pi/1 radian

OpenStudy (hba):

Here is the formula, \[a.b=|a||b|\cos \theta\] Can you solve it now ?

OpenStudy (anonymous):

i tried that,but it seems I am getting the cos teta portion wrong. i calculated as follow 5x2xcos(pi/1) was wrong. and did 5x2x1/2 still wrong

OpenStudy (anonymous):

does it really say \(\frac{\pi}{1}\) ??

OpenStudy (hba):

Cos(pi)=Cos (180)=-1

OpenStudy (anonymous):

Yes, i got it. it is -10 thanks...but how did you derieved at cos(180)...there where I am confused.

OpenStudy (hba):

You wanna know how to convert it from radian to degrees ?

OpenStudy (anonymous):

yes, oh so i have to convert it to degree all the time right?

OpenStudy (anonymous):

i would forget degrees, since you are working in radians one complete rotation is \(2\pi\) , half a rotation is \(\pi\) think only of where you are on the unit circle 

OpenStudy (hba):

Not if you learn the table. But remember \[\pi=180^0\]

OpenStudy (hba):

pi radians*

OpenStudy (anonymous):

@hba will show you the correct way, and i will shut up, but a piece of advice is to forget degrees when you are working in radians and familiarize yourself with the radian measure of angles as corresponds to the unit circle this will make for much less confusion down the line

OpenStudy (anonymous):

that is right,but converting to degree is much simple way to solve the problem...thanks a lot

OpenStudy (hba):

Welcome ;)